OpenStudy (anonymous):

find the standard form of the equation of the parabola with a focus at (-7,0) and a directrix at x=7. PLEASE HELP WILL GIVE MEDALS!!!!

OpenStudy (anonymous):

How do you think your supposed to do this?

OpenStudy (anonymous):

tbh im not sure.

OpenStudy (anonymous):

my options are x=(-1/28)y^2 -28y-=x^2 y^2=-14x y=(-1/28)x^2

OpenStudy (michele_laino):

Hint: we can apply the definition of a parabola

OpenStudy (anonymous):

is there an equation I can use

OpenStudy (michele_laino):

let's suppose that P=(x,y) be a point which belongs to our parabola, then we can write: \[\sqrt {{{\left( {x - \left( { - 7} \right)} \right)}^2} + {{\left( {y - 0} \right)}^2}} = \left| {x - 7} \right|\]

OpenStudy (michele_laino):

now, we have to square both sides of that equation, and solve it for x

OpenStudy (anonymous):

but wouldn't the absolute value part just be a positive and a negative?

OpenStudy (michele_laino):

the absolute part is always positive, since at the left side, we have a square root, which, by definition, is positive.

OpenStudy (anonymous):

that one has the wrong focus, and directrix

OpenStudy (anonymous):

what does the standard form look like?

OpenStudy (michele_laino):

I think that my equation is right. Please try to square both sides, and you will get one of your options.

OpenStudy (anonymous):

okay just give me one sec

OpenStudy (anonymous):

okay its not working my calculator sucks

OpenStudy (michele_laino):

here are your steps:

OpenStudy (michele_laino):

\[\begin{gathered} \sqrt {{{\left( {x - \left( { - 7} \right)} \right)}^2} + {{\left( {y - 0} \right)}^2}} = \left| {x - 7} \right| \hfill \\ {\left( {x + 7} \right)^2} + {y^2} = {\left( {x - 7} \right)^2} \hfill \\ {x^2} + 49 + 14x + {y^2} = {x^2} + 49 - 14x \hfill \\ 28x + {y^2} = 0 \hfill \\ x = - \frac{{{y^2}}}{{28}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

so how does it get to become x=-(1/28)y^2?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

we have to develop the algebraic computation. I have started using the definition of a parabola, namely: "a parabola is that curve whose points have the same distance from the focus and from the directrix".

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

the distance of our generic point P=(x,y) from the focus is: \[\sqrt {{{\left( {x - \left( { - 7} \right)} \right)}^2} + {{\left( {y - 0} \right)}^2}} \]

OpenStudy (michele_laino):

whereas the distance of the generic point P=(x,y) from the directrix, is: \[\left| {x - 7} \right|\]

OpenStudy (anonymous):

yes and then we squared each side right?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so the answer is x=(-1/28)y^2

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thank You so much!

OpenStudy (michele_laino):

Thank You!